OpenStudy (anonymous):

Solve 2sin^2(2x)=1 So far I got: 8sin^2(x)cos^2(x)=1 Using sin(2x)

OpenStudy (anonymous):

\[2\sin ^{^{2}}2x = 1\] I got so far: \[8\sin ^{2}x \times \cos ^{2}x = 1\]

OpenStudy (anonymous):

Having cos(x) throws me off. ...Do I express cos(x) in terms of sin(x)?